OpenStudy (anonymous):

area between two graphs. y=2 and y=x^2-2

OpenStudy (anonymous):

You know how to integrate?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. So to find the area under the function you just integrate it right?

OpenStudy (anonymous):

set the two equations equal to find the bounds

OpenStudy (anonymous):

Since the area under the top function is the area between them and the area under the bottom function. The area between them would be Area under top function - Area under bottom function

OpenStudy (anonymous):

So find the x values for which the two functions are the same, then integrate each function between those bounds, and subtract the resulting areas. Sometimes it's easiest to combine the two integrals and simplify before evaluating the integral, but sometimes it's not.

OpenStudy (anonymous):

would it be 10.66 repeating?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Actually, since your parabola is under the y=0 line for a bit this method won't quite work. You'll need to break up the integrals. You'll have \[Area = \int\limits_{-2}^{-\sqrt{2}}[2 - (x^2-2)]dx - \int\limits_{-\sqrt{2}}^{\sqrt{2}}(x^2-2)dx + \int\limits_{\sqrt{2}}^{2}[2 - (x^2-2)]\]

OpenStudy (anonymous):

thats nonsence... just integrate \[\int\limits\limits_{-2}^{2}2-x ^{2}+2 dx=\int\limits\limits_{-2}^{2}4-x ^{2} dx\]

OpenStudy (anonymous):

That doesn't work. Because part of that integral will be negative area.

OpenStudy (anonymous):

no, since you are adding 2 to the parabola in fact. it never crosses x-axis

OpenStudy (amistre64):

for practice, let me see what I can mess up :) I agree that the interval is [-2,2], but since we have mirror images across the yaxis, we can just do [0,2] and double it. y = 2 is just a box that is 2(2 by 2) = 8 (S)x^2-2 dx -> x^3/3 -2x |0,2 = 2[(8/3) -6] = 16/3 -36/3 = -20/3 8 --20/3 = 24/3 + 20/3 = 44/3 I would say, and this is my guess.. that the area between the 2 curves is 44/3. Now just let me know what I did wrong :)

OpenStudy (anonymous):

everything :) \[x^{3}/3 - 2x \] from 0 to 2 is 8/3-4=8/3-12/3=-4/3 *2 => -8/3 8-(-8/3)=32/3=10+2/3=10.66666.... :D